OpenStudy (anonymous):

how do you evaluate the equation: y=(x-1)(x+2) Please explain steps, thank you so much for your help!

OpenStudy (anonymous):

y = (a-b)(c+d) = ac + ad -bc -bd

OpenStudy (anonymous):

where do you get a b and c and d?

OpenStudy (anonymous):

those are variables, i'm giving you a general solution.

OpenStudy (anonymous):

ohh so how do you solve that with that im sorry im confused

OpenStudy (anonymous):

y = (x-1)(x+2) = (x)(x) + 2x - x - 2

OpenStudy (anonymous):

y=(x-1)(x+2) Multiply each term in the first group by each term in the second group using the FOIL method. FOIL stands for First Outer Inner Last, and is a method of multiplying two binomials. First, multiply the first two terms in each binomial group. Next, multiply the outer terms in each group, followed by the inner terms. Finally, multiply the last two terms in each group. y=(x*x+x*2-1*x-1*2) Multiply x by x to get x^(2). y=(x^(2)+x*2-1*x-1*2) Multiply x by 2 to get 2x. y=(x^(2)+2x-1*x-1*2) Multiply -1 by x to get -x. y=(x^(2)+2x-x-1*2) Multiply -1 by 2 to get -2. y=(x^(2)+2x-x-2) According to the distributive property, for any numbers a, b, and c, a(b+c)=ab+ac and (b+c)a=ba+ca. Here, x is a factor of both 2x and -x. y=(x^(2)+(2-1)x-2) To add integers with different signs, subtract their absolute values and give the result the same sign as the integer with the greater absolute value. In this example, subtract the absolute values of 2 and -1 and give the result the same sign as the integer with the greater absolute value. y=(x^2+(1)x-2) Remove the parentheses. y=(x^2+x-2) Remove the parentheses around the expression x^(2)+x-2. y=x^2+x-2

OpenStudy (anonymous):

THankk you so much!

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Can I ask you another question so I foiled it and got x*2+x-2 how do i do the next step?

OpenStudy (anonymous):

Are you trying to solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh, then it's easier than all the steps I wrote down.

OpenStudy (anonymous):

Here. y=(x-1)(x+2) set (x-1)=0 and(x+2)=0 Then solve each equation x-1=0......x=1 x+2=0......x=-2

OpenStudy (anonymous):

Or you could solve: y=x^2+x-2, by using the quadratic formula.

OpenStudy (anonymous):

thank you so much for your help !